a rectangular room is 1.8 times as long as it is wide, its perimeter is 29 meters. find the diminsion of the room
Found 2 solutions by Boreal, mananth:
Answer by Boreal(15235)

perimeter is 2L+2W, and half-perimeter is L+W=14.5 m
L=1.8W, so the half-perimeter is 2.8W=14.5m
W=5.18 m
L=1.8*5.18=9.32 m
2W=10.36 m
2L=18.64 m
that sum is 29 m, so it checks.
5.18 m x 9.32 m

Answer by mananth(16946)

Perimeter = 29
Length =1.8 x
Width =x ft

Perimeter = 2 1.8 x +x
Perimeter = 2 1.8 x+ 1 x
Perimeter = 3.6 x + + 2 x
29 = 5.6 x
29 = 5.6 x
29 = 5.6 x

x= 5.2 ft Width
Length =9.3ft